OpenStudy (anonymous):

A certain field is a rectangle with a perimeter of 820ft. The length is 188ft more than the width. Find the length and width of the rectangular field?

OpenStudy (anonymous):

2l+2w=820 l+w=410 l=w+188 w+w+188=410 2w+188=410 w+94=205 w=111 l=111+188 l=299

OpenStudy (anonymous):

thanks